Find the indicated complement A certain group of women has a 0.15% rate of red/green color blindness. If a woman is randomly selected, what is the probability that she does not have red/green color blindness? What is the probability that the woman selected does not have redigreen color blindness? Type an integer or a decimal. Do not round.)
Answer #1

Given that a certain group of women has a 0.15% rate of red/green color blindness.

Thus, P(A)=0.15%=0.15/100=0.0015

We have to find that if a women is randomly selected,what will be the probability that she does not have red/green color blindness. Thus we have to find P(A').

We know, P(A)+P(A')=1

or, P(A')=1-P(A)=1-0.0015=0.9985

Thus the required probability is 0.9985.